I am having difficulty solving this problem, its a little more complicated than the last. Please help.
5(y+1)+2=8(y-1)+3

Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

5(y+1)+2=8(y-1)+3
5y+5+2 = 8y-8+3
-----
5y+7 = 8y-5
3y = 12
y = 4
---------------
Cheers,
Stan H.

Answer by ewatrrr(24785)


Hi,
using the distributive Principle and combining Like terms
5(y+1)+2=8(y-1)+3
| Applying distibutive Principle
5y +5 + 2 = 8y - 8 + 3
5y + 7 = 8y - 5
| combining like Terms by subtracting 5y from both sides of the equation
7 = 3y - 5
| combining like Terms by adding 5 to both sides of the equation
12 = 3y
| isolating y, the variable, on one side of the equation
| Divided both sides by 3
4 = y
CHECKING our Answer***
5(y+1)+2=8(y-1)+3
5*5 + 2 = 8*3 + 3 = 27